a)solve 5x^2=10
b)find the x intercepts of f(x)=5x^2-10
^means it is raising to that power

Answer by ewatrrr(24785)

 
Hi,
a)solve 5x^2=10 x = ± sqrt(2)
b)find the x intercepts of f(x)=5x^2-10
f(x)=5x^2-10
f(x)=5(x^2-2)
x intercepts (when f(x) = 0)
0 = 5(x^2-2)
2 = x^2
x = ± sqrt(2)
y-intercepts are ( sqrt(2), 0) and (-sqrt(2),0)